[Hallc_running] [New Logentry] RC Daily Update

pooser at jlab.org pooser at jlab.org
Wed Nov 14 14:35:01 EST 2018

Logentry Text:
Beam was restored to the hall at the start of swing shift yesterday evening and data taking on kinematic setting 6-2 continued.  At approximately 17:35 we switched to kinematic setting 6-4 (https://logbooks.jlab.org/entry/3625702).

The start of run auto-logs were missing for runs 6176 and 6177 (https://logbooks.jlab.org/entry/3625763) however, it was recovered by a restart of the DAQ. At ~22:35 we switched to kinematic setting 6-9 (https://logbooks.jlab.org/entry/3625810) and then switched to kinematic group 6-11 at 23:12 (https://logbooks.jlab.org/entry/3625833).  

At 23:41 Q2 initiated a slow discharge and threw an interlock however, an alarm was never initiated (https://logbooks.jlab.org/entry/3625844).  In order to recover Q2 an access was required to perform a hard reset on the PLC for Q2 (https://logbooks.jlab.org/entry/3625862, https://logbooks.jlab.org/entry/3625900).  Beam was restored to the hall at ~00:45 and data taking on kinematic group 6-11.

Data taking remained smooth for the majority of the night until the hall, BSY and injector dropped to power permit (https://logbooks.jlab.org/entry/3625978) around 07:00 and was likely attributed to the SHMS door not being shut entirely and causing a PSS trip.  It was the B chain on the SHMS run/safe box dropped which dropped the hall to power permit (https://logbooks.jlab.org/entry/3625990).  MCC had to then cycle the BSY and injector in order to bring up the other halls.

Beam was restored in the hall at ~09:10 and data taken on kinematic group 10-3 has commenced.  Moving forward we will continue data taking with kinematic group 10-4 and following the posted run plan.

Notices in effect for shift crews:
Split the large desired charge between multiple runs (~1 hour per run)
Listed currents are estimates. For runs at less than 60uA, increase current until SHMS singles rates are between 400k-500k.
How-to power cycle PLC I/O chassis: https://logbooks.jlab.org/entry/3625862, https://logbooks.jlab.org/entry/3625900

Planned accesses to the hall:

Planned accelerator configuration changes:
Monday, Nov 19: system recovery and spot move, estimated 4 hr downtime although it is scheduled for 8.
Tuesday, Nov 20: CHL UPS battery change, should be noninvasive however, induce CHL anomaly is possible
Wednesday, Nov 21: Hall A goes offline and MCC wants to conduct high current beam study of energy spread at 40 uA.  It is unclear that we will be effected
Monday, Nov 26: Energy change to 1.86 GeV/pass, we will be moved over to 500 MHz laser.

Special requests:
Carlos Munoz Camacho would like access to the hall to show Hall-C to two French engineers  working on the NPS project.  It should be a ~30 minute access, and is a low priority.  He can be contacted at 357-1065

We will meet today at 4:30 pm in the second floor of the counting house.  

* Bluejeans remote connection information:
  To join the Meeting:
  To join via Room System:
  Video Conferencing System: bjn.vc -or-
  Meeting ID : 861439920
  To join via phone :
  1)  Dial:
    +1.408.740.7256 (US (San Jose))
    +1.888.240.2560 (US Toll Free))
    +1.408.317.9253 (US (Primary, San Jose))
    (see all numbers - http://bluejeans.com/numbers)
  2)  Enter Conference ID : 861439920


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3626178
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20181114/7b236f9d/attachment.html>

More information about the Hallc_running mailing list